OpenStudy (anonymous):

Identify the vertex and the axis of symmetry of the graph of the function y = 3(x + 2)2 – 3. vertex: (2, –3); axis of symmetry: x = 2 vertex: (–2, –3); axis of symmetry: x = –2 vertex: (2, 3); axis of symmetry: x = 2 vertex: (–2, 3); axis of symmetry: x = –2

OpenStudy (anonymous):

this is a vertex form : y=a(x-h)^2 +k with vertex (h,k)

OpenStudy (anonymous):

hence the axis of symmetry is x =h

OpenStudy (anonymous):

can you answer now?

OpenStudy (anonymous):

I dont understand how to write it ?

OpenStudy (anonymous):

general form : y=a(x-h)^2 +k our case : y = 3(x + 2)2 – 3 so for us a = 3 h = -2 k = -3

OpenStudy (anonymous):

is the vertex ( 2, -3 )

OpenStudy (anonymous):

try again using what i just wrote

OpenStudy (anonymous):

I dont understand what x is ?

OpenStudy (anonymous):

x is the independent variable .. given y=a(x-h)^2 +k the vertex is (h,k) and axis of symmetry is x =h

OpenStudy (anonymous):

sorry my internet connection is bad

OpenStudy (anonymous):

im confused :/